OpenStudy (anonymous):

Lois wrote the equation shown below. 4x + 12y = 4(x + 3y) Which property did Lois apply?

OpenStudy (anonymous):

commutative property of addition associative property of addition distributive property of addition addition property of equality

Parth (parthkohli):

That dude is smart. He was correct. We usually have questions like "At what step was Lois wrong?" :P

Parth (parthkohli):

Well. Do you have any guesses for this one?

OpenStudy (anonymous):

ummm... b?

OpenStudy (anonymous):

Something to help out you in general with these type questions: apply process of elimination first! :-)

Parth (parthkohli):

\( \color{Black}{\Rightarrow a + (b + c) = (a + b) + c}\) Associative ^ \( \color{Black}{\Rightarrow a( b + c) = ab + ac }\) Distributive ^ \( \color{Black}{\Rightarrow a + b = b + a }\) Commutative ^

Parth (parthkohli):

Which one in the examples according to you is closest?

OpenStudy (anonymous):

the 2nd

Parth (parthkohli):

I don't think so.

OpenStudy (anonymous):

??

Parth (parthkohli):

2nd one doesn't look like the property that Lois was talking about.

OpenStudy (anonymous):

oooooo commutative

Parth (parthkohli):

Nah

OpenStudy (anonymous):

what ? thn im lost

Parth (parthkohli):

How did you get commutative and associative right there? Can you explain?

Parth (parthkohli):

Hint: look at the thing as \(4(x + 3y) = 4x + 12y\) instead.

OpenStudy (anonymous):

so it is assos

Parth (parthkohli):

How so? If you think that it is associative, then can you tell me what a, b and c are in this example?

OpenStudy (anonymous):

because it looks the same as the example

Parth (parthkohli):

Well. Can you tell me what \(a,b,c\) are in the example?

Parth (parthkohli):

The associative property says \(a + (b + c) = (a + b) + c\).

OpenStudy (anonymous):

thats wat im gonna go with

Parth (parthkohli):

What are a,b and c?

OpenStudy (anonymous):

commutative property of addition associative property of addition distributive property of addition

Parth (parthkohli):

Not the choices..

OpenStudy (anonymous):

wat do u think it is

Parth (parthkohli):

It's one of the choices.

OpenStudy (anonymous):

Its distributive... which is what she prolly meant when she said the second one^^